# Variational Quantum Framework for Partial Differential Equation Constrained Optimization

### v1

@@ 4.1.3 Ansatz Selection
The VQLS employs a parameterized quantum circuit (PQC) or an ansatz which is a gate sequence $\mathbf{V}(\mathbf{\theta})$ which simulates a potential solution $|\mathbf{\psi}(\mathbf{\theta})\rangle=\mathbf{V}(\mathbf{\theta})|\mathbf{0}\rangle$. There are variety of choices for ansatz which broadly fall into the categories of fixed-structure ansatz, problem-specific ansatz, and dynamic/adaptive ansatz [9, 33]. One popular choice of fixed-structure ansatz is the hardware-efficient ansatz (HEA). HEA are designed without taking into account the specific problem being solved, but rather only the topology (backend connectivity of the qubits) and available gates of a specific quantum computer. These ansatz can be constructed to be more resistant to noise on any specific available quantum device, but they may fall short of finding a solution, as any particular HEA is not guaranteed to span the region of the Hilbert space containing any good approximation of the solution. Furthermore, HEA can suffer from barren plateau resulting in trainability issue for large scale problems.
Problem-specific ansatz does not take into account the specific quantum device being used, and rather tries to exploit the knowledge of the problem available. The Quantum Alternating Operator Ansatz (QAOA) [37] is one such proposed problem-specific ansatz, using two Hamiltonians, known as the driver and the mixer, constructed from specific knowledge of the problem. The requirement of Hamiltonian simulation from the QAOA makes these anatz far less near-term. In variable structure or dynamic ansatz one optimizes over both the gate parameters $\mathbf{\theta}$ and the gate placement in the circuit [33].
For real valued problems, such as the example heat transfer optimization problem, the quantum gates used in the ansatz can be restricted appropriately. In this work, we use the fixed structure real-valued ansatz shown in Fig. 3. This is the modified circuit 9 from [38] that was the least expressible with a high KL divergence value of 0.68.
Figure 3: Real-valued ansatz used in this study.
@@ 4.1.4 Cost Function
with $|0_{k}\rangle$ being a zero state on qubit $k$, and $\mathbf{I}_{\tilde{k}}$ being identity on all qubits except the qubit $k$. The normalized local cost function can be expressed as
$C_{l}(\mathbf{\theta})=1-\frac{1}{2n}\frac{\sum_{i=1}^{n_{l}}\sum_{j=1}^{n_{l}}\sum_{k=1}^{n}\delta_{ijk}\alpha_{i}\alpha_{j}^{*}}{\langle\mathbf{\phi}|\mathbf{\phi}\rangle},$
where, the expression for $\langle\mathbf{\phi}|\mathbf{\phi}\rangle$ is as given above, and
$\delta_{ijk}$ $=$ $\langle\mathbf{0}|\mathbf{V}^{*}\mathbf{A}_{j}^{*}\mathbf{U}_{b}(|0_{k}\rangle\langle 0_{k}|\otimes\mathbf{I}_{\tilde{k}})(\mathbf{U}_{b})^{*}\mathbf{A}_{i}\mathbf{V}|\mathbf{0}\rangle,$ (68)
$=$ $\beta_{ij}+\langle\mathbf{0}|\mathbf{V}^{*}\mathbf{A}_{j}^{*}\mathbf{U}_{b}(\mathbf{Z}_{k}\otimes\mathbf{I}_{\tilde{k}})(\mathbf{U}_{b})^{*}\mathbf{A}_{i}\mathbf{V}|\mathbf{0}\rangle.$
Note that to calculate $\delta_{ijk}$ one only needs to measure the real/imaginart parts of the $\langle\mathbf{0}|\mathbf{V}^{*}\mathbf{A}_{j}^{*}\mathbf{U}_{b}(\mathbf{Z}_{k}\otimes\mathbf{I}_{\tilde{k}})(\mathbf{U}_{b})^{*}\mathbf{A}_{i}\mathbf{V}|\mathbf{0}\rangle$ which can be accomplished via a slight variation of Hadamard test circuit, see [9] for the details.
@@
where, $\epsilon=\rho(\psi,\psi(\mathbf{\theta}_{*}))$ is the error in the normalized solution with $|\psi\rangle$ being the exact solution of Eqn. 59, and $|\psi(\mathbf{\theta}_{*})\rangle$ being the approximate VQLS solution. These relations can be used in convergence criterion as discussed later.
@@ 4.1.6 VQLS Algorithm
The steps for VQLS are summarized in Algorithm 1. To initialize the VQLS iterations (see line (3)) the parameter vector $\mathbf{\theta}_{in}$ needs to be prescribed. Randomly choosing these parameters with highly likelihood will cause VQLS to start in the region of barren plateau which is not desirable. To overcome these, several other methods for initialization can be employed, see for example [40, 41]. From the relations (69), one can select threshold $\gamma$ such that, for instance, for $C_{l}\leq\gamma$, results in approximation error $\epsilon\leq\kappa\sqrt{\gamma\log N}$. The number of shots $n_{sh}$ determines the number of times the quantum circuit is run to estimate each of the term (i.e., $\beta_{ij},\gamma_{ij},\delta_{ijk}$, see the Sec. 4.1.4) in the cost evaluation. Finally, any of the classical optimizers discussed in the previous section can be used in the Step 6.
@@
3:Initialize $k=0$, $\mathbf{\theta}^{0}=\mathbf{\theta}_{in}$ and $c_{l}=\infty$.
4:while $c_{l}\geq\gamma$ do
5:   Compute cost $c_{l}=C_{l}(\mathbf{\theta}^{k})$ using associated quantum circuits with $n_{sh}$ number of shots.
@@ 4.2.2 Acquisition Function
where, $f^{*}=\min\{y_{i}|y_{i}\in\mathcal{D}\}$ and $\mu(\mathbf{p})$ and $\sigma(\mathbf{p})$ are as defined in Eqn. 71 with $\Sigma_{\epsilon}=0$. This expectation has a closed form expression
@@ 4.2.3 BO Algorithm
5:Update the GP posterior mean and covariance using Eqn. 71, with $\Sigma_{\epsilon}=\mbox{diag}(\Delta_{1},\cdots,\Delta_{N_{o}})$.
@@
5 Bi-Level Variational Quantum PDE Constrained Optimization (BVQPCO) Framework
$g_{i}(\mathbf{p})\leq 0,i=1,\cdots,N_{e},$
where, $\mathbf{p}\in R^{N_{p}}$ are the design variables, $\mathbf{\psi}\in\mathbb{R}^{n_{s}}$ is the discretized solution vector of the underlying PDE represented in the form of the linear constraints (76), $g_{i}(\mathbf{p})$ are the constraints on $\mathbf{p}$, and $C_{d}$ is the cost function which can expressed as a function $f$ involving $\mathbf{\phi}_{d},\mathbf{\phi}_{n}\in\mathbb{R}^{n_{s}}$ as fixed vectors associated with the cost function.
Figure 4: Flow diagram of the BVQPCO framework. VQLS is used for inner level optimization to solve the linear system constraints, arising from the discretization of the underlying PDEs, for given design parameters, and evaluate the quantities related to design cost/objective function. The black box optimizer is used in the outer level to select next set of parameters values based on the evaluated design cost.
To solve this problem in a variational quantum framework, we propose BVQPCO framework, where the outer optimization level iteratively selects $\mathbf{p}^{k}$ based on the solution $|\mathbf{\psi}\rangle$ obtained via VQLS for the linear system $\mathbf{A}(\mathbf{p}^{k})|\mathbf{\psi}\rangle=|\mathbf{b}(\mathbf{p}^{k})\rangle$. Thus, the optimization problem (74) can be expressed in quantum form as,
$g_{i}(\mathbf{p})\leq 0,i=1,\cdots,N_{e},$
where, and $|\mathbf{\phi}_{d}\rangle,|\mathbf{\phi}_{n}\rangle\in R^{n_{s}}$ define quantum analogue of the cost function parameters. Note that since,
the cost function (74) can be equivalently expressed in the quantum form (77). Fig. 4 shows the overall flow diagram of the BVQPCO framework.
The steps to solve (77) are described in Algorithm 3. Some remarks follow:
  • Depending on how $\mathbf{A}$ depends on the parameters $\mathbf{p}$, a parameter dependent LCU $\{\alpha_{i}(\mathbf{p}),\mathbf{A}_{i}\}$ can be pre computed once, thus saving computational effort. For instance, for the optimization problem (57), $\mathbf{A}$ for both the explicit (see Eqn. 36) and implicit (see Eqn. 55) discretization schemes has a separable structure. Thus, LCU for the matrices $\tilde{\mathbf{A}}_{f1},\tilde{\mathbf{A}}_{f2}$ (and similarly $\tilde{\mathbf{A}}_{b1},\tilde{\mathbf{A}}_{b2}$) can be computed once, and the LCU of $\mathbf{A}$ can be readily constructed for any choice of parameters $\mathbf{p}=(l,\alpha)$.
  • For the inner optimization for VQLS in the line (7), one can use any of the optimizer as discussed in Sec. 4.1.5.
  • For the outer optimization in line (8) one can use any black box optimization method. We propose to use BO as described in Sec. 4.2.
2:Output: Optimal parameters: $\mathbf{p}^{*}$
3:Initialize $k=0$ and $\mathbf{p}^{0}=\mathbf{p}_{in}$.
5:while stopping criteria not met do
6:   Compute LCU $\{\alpha_{i}(\mathbf{p}^{k}),\mathbf{A}_{i}(\mathbf{p}^{k})\}_{i=1}^{n_{k}}$ and unitary $\mathbf{U}_{b}^{k}$ such that $|\mathbf{b}(\mathbf{p}^{k})\rangle=\mathbf{U}_{b}^{k}|\mathbf{0}\rangle$.
7:   Compute $\mathbf{\theta}_{*}^{k}=VQLS(\{\alpha_{i}(\mathbf{p}^{k}),\mathbf{A}_{i}(\mathbf{p}^{k})\}_{i=1}^{n_{k}},\mathbf{U}_{b}^{k},\mathbf{V}(\mathbf{\theta}),\gamma,n_{sh})$.
9:   Use classical optimizer to select next $\mathbf{p}^{k+1}$ subject to constraints $g_{i}(\mathbf{p}^{k+1})\leq 0,i=0,\cdots,N_{e}$.
11:   $k\leftarrow k+1$
12:end while
13:Return $\mathbf{p}^{k}$
Algorithm 3 Bi-level Variational Quantum Optimization Algorithm.
6 Error and Computational Complexity Analysis
@@
with $\rho$ being the trace norm, $|\psi\rangle$ being the exact solution of Eqn. 59, and $|\psi(\mathbf{\theta}_{*})\rangle$ being the approximate VQLS solution.
@@ 6.1 Error Analysis for VQLS Based Explicit Euler Linear ODE Solver
Applying the forward Euler method to the system with step size $h$ yields
@@
The optimal parameter $\mathbf{\theta}_{*}$ can be used to prepare a solution $|\overline{\mathbf{u}}_{f}\rangle=\mathbf{V}(\mathbf{\theta}_{*})|\mathbf{0}\rangle$ which is an approximation to $|\tilde{\mathbf{u}}_{f}\rangle$. The approximation error is characterized by the Lem. 1.
@@
where, $N,M>1$. Let the VQLS algorithm be terminated under the condition
$C_{l}\leq\gamma,$
@@ 6.2 Error Analysis for VQLS Based Implicit Euler Linear ODE Solver
Let $h$ be chosen such that $h\nu\leq 0.5$, where $\nu=\|\mathbf{A}\|$. Then the stiffness $\kappa$ of $\tilde{\mathbf{A}}_{b}$ in ( 104) is bounded as follows
@@ 7 Numerical Results
In this section, we demonstrate the application of our BVQPCO framework to solve the heat transfer optimization problem described in  Sec. 3.1. The problem is discretized on a grid of size $N=8$ and $\Delta x=\frac{l}{N-1}=1/7$. We consider the implicit time discretization given in Sec. 3.1.2 with $\Delta t=0.25s$ and $M=4$ time steps. The following values of the parameters are used: $q_{0}=50.0,k=1.0,l_{\min}=2.0,l_{\max}=4.0,\alpha_{\min}=0.2,\alpha_{\max}=0.3,w_{1}=10.0,w_{2}=1.0,w_{3}=5.0$. There is a gradient in the initial temperature distribution across the grid as shown in Fig. 5 at $t=0s$.
@@
The implementation of Bayesian optimization from Ax, BoTorch [46] was used with Ax’s Loop API. Sobol sequences are used to generate ten initial samples in the parameter space and twenty BO iterations are performed. Fixed noise Gaussian process (where noise is provided as an input) is used as the surrogate model and qNoisy EI is used as the acquisition function. While we do not consider device or measurement noise in this work, we use qNoisy EI model to capture the error in estimation of design cost from the inner VQLS optimization loop. We input standard error to the BO model as $\sigma_{\epsilon}(l,\alpha)\approx c\sqrt{C_{g}(l,\alpha)\log N}$ using Eqn. 69, where $c$ is a fixed constant we selected empirically.
@@ 7.1 Convergence of VQLS
7.2 Convergence of the bi-level framework
@@ 8 Conclusion
In this paper we presented a novel variational quantum framework for PDE constrained design optimization problems. The proposed framework utilizes the VQLS algorithm and a black box optimizer nested in bi-level optimization structure. We presented a detailed complexity analysis to establish potential advantage of our framework over classical techniques. We demonstrated the framework on a heat transfer optimization problem and presented simulation results using Bayesian optimization as the black box optimizer. The analysis and results demonstrate the correctness of our framework for solving simulation-based design optimization problems.
Future work involves extending our bi-level optimization framework to nonlinear PDE constrained problems by leveraging our prior work on Carleman linearization based embedding approach [49]. One of the bottlenecks in the proposed framework is the computation of VQLS cost functions which scales polynomially with the number of LCU terms. We can employ a more efficient tensor product decomposition [50], to provide exponential speed-up in the VQLS cost evaluation by exploiting the underlying structure and sparsity of matrices such as arising from PDE discretizations. An adaptive multigrid strategy to accelerate convergence of variational algorithms was recently proposed in [51]. Such techniques can be adapted in the context of VQLS to further improve our framework and scale to larger problem sizes.

### v2

@@ 4.2.3 Ansatz Selection
The VQLS employs a parameterized quantum circuit (PQC) or an ansatz which is a gate sequence $\mathbf{V}(\mathbf{\theta})$ which simulates a potential solution $|\mathbf{\psi}(\mathbf{\theta})\rangle=\mathbf{V}(\mathbf{\theta})|\mathbf{0}\rangle$. There are variety of choices for ansatz which broadly fall into the categories of fixed-structure ansatz, problem-specific ansatz, and dynamic/adaptive ansatz [9, 48]. One popular choice of fixed-structure ansatz is the hardware-efficient ansatz (HEA). HEA are designed without taking into account the specific problem being solved, but rather only the topology (backend connectivity of the qubits) and available gates of a specific quantum computer. While these ansatz can be constructed to be more resistant to noise on any specific available quantum device,, they may not be sufficiently expressive and suffer from barren plateau resulting in trainability issue for large scale problems. Problem-specific ansatz does not take into account the specific quantum device being used, and rather tries to exploit the knowledge of the problem available. The Quantum Alternating Operator Ansatz (QAOA) [7] is one such proposed problem-specific ansatz, using two Hamiltonians, known as the driver and the mixer, constructed from specific knowledge of the problem. The requirement of Hamiltonian simulation from the QAOA makes these anatz far less near-term. In variable structure or dynamic ansatz one optimizes over both the gate parameters $\mathbf{\theta}$ and the gate placement in the circuit [48].
For real valued problems, such as the example heat transfer optimization problem in Sec. 3.1, the quantum gates used in the ansatz can be restricted appropriately. In this work, we use the fixed structure real-valued ansatz shown in Fig. 4. This is the modified circuit 9 from [52] that was the least expressible with a high KL divergence value of 0.68.
Figure 4: Real-valued ansatz (modified circuit 9) on 4 qubits comprising of Hadamard gate (H), SWAP gates and rotation about y-axis gates (RY).
@@ 4.2.4 Cost Function
with $|0_{k}\rangle$ being a zero state on qubit $k$, and $\mathbf{I}_{\tilde{k}}$ being identity on all qubits except the qubit $k$. Computation of $C_{l}$ can be accomplished via a slight variation of Hadamard test circuit, see [9] for the details.
@@
where, $\epsilon=\rho(\psi,\psi(\mathbf{\theta}_{*}))$ is the error in the normalized solution with $|\psi\rangle$ being the exact solution of Eqn. 66, and $|\psi(\mathbf{\theta}_{*})\rangle$ being the approximate VQLS solution. These relations can be used in convergence criterion as discussed later.
@@ 4.2.6 VQLS Algorithm
The steps for VQLS are summarized in Algorithm 2. To initialize the VQLS iterations (see line (3)) the parameter vector $\mathbf{\theta}_{in}$ needs to be prescribed. Randomly choosing these parameters with highly likelihood will cause VQLS to start in the region of barren plateau which is not desirable. To overcome these, several other methods for initialization can be employed, see for example [54, 55]. From the relations (75), one can select threshold $\gamma$ such that, for instance, for $C_{g}\leq\gamma$, results in approximation error $\epsilon\leq\kappa\sqrt{\gamma\log N}$. The number of shots $n_{sh}$ determines the number of times the quantum circuit is run to estimate each of the term (i.e., $\beta_{ij},\gamma_{ij}$, see the Sec. 4.2.4) in the cost evaluation. Finally, any of the classical optimizers discussed in the previous section can be used in the Step 6.
@@
3:Initialize $k=0$, $\mathbf{\theta}^{0}=\mathbf{\theta}_{in}$ and $c=\infty$.
4:while $c\geq\gamma$ do
5:   Compute cost $c=C_{g}(\mathbf{\theta}^{k})$ using associated quantum circuits with $n_{sh}$ number of shots.
@@ 4.3.2 Acquisition Function
where, $f^{*}=\min\{y_{i}|y_{i}\in\mathcal{D}\}$ and $\mu(\mathbf{p})$ and $\sigma(\mathbf{p})$ are as defined in Eqn. 77 with $\Sigma_{\epsilon}=0$. This expectation has a closed form expression
@@ 4.3.3 BO Algorithm
5:Update the GP posterior mean and covariance using Eqn. 77, with $\Sigma_{\epsilon}=\mbox{diag}(\Delta_{1},\cdots,\Delta_{N_{o}})$.
@@
5 Computational Error and Complexity Analysis
@@
with $\rho$ being the trace norm, $|\psi\rangle$ being the exact solution of Eqn. 66, and $|\psi(\mathbf{\theta}_{*})\rangle$ being the approximate VQLS solution.
@@ 5.1 Error Analysis for VQLS Based Explicit Euler Linear ODE Solver
Applying the forward Euler method to the system with step size $h$ yields the difference equation
@@
The optimal parameter $\mathbf{\theta}_{*}$ can be used to prepare a solution $|\overline{\mathbf{u}}_{f}\rangle=\mathbf{V}(\mathbf{\theta}_{*})|\mathbf{0}\rangle$ which is an approximation to $|\tilde{\mathbf{u}}_{f}\rangle$. The approximation error is characterized by the Lem. 1Fig. 5 shows a schematic of the steps involved in transforming and solving a given PDE using VQLS and the associated variables/approximation errors.
ODE $\dot{\mathbf{u}}$ $=$ $\mathbf{A}\mathbf{u}+\mathbf{b}(t),$ $\mathbf{u}(0)$ $=$ $\mathbf{u}_{0},$ $\mathbf{u}_{c}=\left(\begin{array}[]{cc}\mathbf{u}(0)\\ \mathbf{u}(h)\\ \vdots\\ \mathbf{u}(h(M-1))\end{array}\right).$ Difference Equation $\hat{\mathbf{u}}_{f}^{k+1}$ $=(\mathbf{I}+h\mathbf{A})\hat{\mathbf{u}}_{f}^{k}+h\mathbf{b}^{k},$ $\hat{\mathbf{u}}_{f}^{1}$ $=\mathbf{u}(0),$ $\tilde{\mathbf{u}}_{f}$ $=\left(\begin{array}[]{c}\hat{\mathbf{u}}_{f}^{1}\\ \hat{\mathbf{u}}_{f}^{2}\\ \vdots\\ \hat{\mathbf{u}}_{f}^{M}\\ \end{array}\right).$ VQLS $\tilde{\mathbf{A}}_{f}\mathbf{V}(\mathbf{\theta})|\mathbf{0}\rangle=|\tilde{\mathbf{b}}\rangle,$ $|\overline{\mathbf{u}}_{f}\rangle=\mathbf{V}(\mathbf{\theta}_{*})|\mathbf{0}\rangle$. PDESpatialDiscretizationEuler DiscretizationLinear System$\||\mathbf{u}_{c}\rangle-|\tilde{\mathbf{u}}_{f}\rangle\|$$\||\tilde{\mathbf{u}}_{f}\rangle-|\overline{\mathbf{u}}_{f}\rangle\|$VQLSSolution
Figure 5: Schematic showing different stages in going from the PDE to the VQLS solution, associated variables and the error in approximations. The case of explicit/forward Euler scheme is shown as an example. Similar flow applies for the case of implicit/backward Euler scheme.
@@
where, $N,M>1$. Let the VQLS algorithm be terminated under the condition (with similar analysis for other VQLS cost functions)
$C_{g}\leq\gamma,$
@@ 5.2 Error Analysis for VQLS Based Implicit Euler Linear ODE Solver
Let $h$ be chosen such that $h\nu\leq 0.5$, where $\nu=\|\mathbf{A}\|$. Then the stiffness $\kappa$ of $\tilde{\mathbf{A}}_{b}$ in ( 103) is bounded as follows
@@ 6 Numerical Results
In this section, we demonstrate the application of our BVQPCO framework to solve the heat transfer optimization problem described in  Sec. 3.1. The problem is discretized on a grid of size $N=8$ and $\Delta x=\frac{l}{N-1}=1/7$. We consider the implicit time discretization given in Sec. 3.1.2 with $\Delta t=0.25s$ and $M=4$ time steps. The following values of the parameters are used: $q_{0}=50.0,k=1.0,l_{\min}=2.0,l_{\max}=4.0,\alpha_{\min}=0.2,\alpha_{\max}=0.3,w_{1}=10.0,w_{2}=1.0,w_{3}=5.0$. There is a gradient in the initial temperature distribution across the grid as shown in Fig. 6 at $t=0s$. This can be achieved by choosing the rotation angles in the state preparation circuit shown in Fig. 3.
@@
The implementation of Bayesian optimization from Ax, BoTorch [60] was used with Ax’s Loop API. Sobol sequences are used to generate ten initial samples in the parameter space and twenty BO iterations are performed. Fixed noise Gaussian process (where noise is provided as an input) is used as the surrogate model and qNoisy EI is used as the acquisition function. While we do not consider device or measurement noise in this work, we use qNoisy EI model to capture the error in estimation of design cost from the inner VQLS optimization loop. We input standard error to the BO model as $\sigma_{\epsilon}(l,\alpha)\approx c\sqrt{C_{g}(l,\alpha)\log N}$ using Eqn. 75, where $c$ is a fixed constant we selected empirically.
@@ 6.1 Convergence of VQLS
6.2 Convergence of the BVQPCO framework
@@ 7 Conclusion
In this paper we presented BVQPCO, a novel variational quantum framework for PDE constrained design optimization problems. The proposed framework utilizes the VQLS algorithm and a black box optimizer nested in bi-level optimization structure. We presented a detailed computational error and complexity analysis to establish potential benefits of our framework over classical techniques. We demonstrated the framework on a heat transfer optimization problem and presented simulation results using Bayesian optimization as the black box optimizer. The analysis and results demonstrate the correctness of our framework for solving simulation-based design optimization problems. To the best of our knowledge, this is the first work of its kind in designing a bi-level variational quantum framework for solving PDE constrained design optimization problems.
Future work will involve studying and mitigating the effect of device/ measurement noise and making the BVQPCO framework robust. It will also be worthwhile to study the scalability of the framework by applying it to larger problem sizes and implementing on quantum hardware. One of the bottlenecks in the BVQPCO framework is the computation of VQLS cost function which scales polynomially with the number of LCU terms. We can employ a more efficient tensor product decomposition [63], which exploits the underlying structure and sparsity of matrices such as arising from PDE discretizations, to provide an exponential speed-up in the VQLS cost evaluation. An adaptive multigrid strategy to accelerate convergence of variational quantum algorithms was recently proposed in [64]. Such techniques can be adapted in the context of VQLS to further improve our framework and scale to larger problem sizes. By leveraging Carleman linearization based embedding approach [26, 28] we also plan to extend our framework for nonlinear PDE constrained optimization problems. Finally, adapting the proposed framework for fault tolerant quantum computing setting is also an avenue for future research.